OpenStudy (anonymous):

an airplane is sighted at angles of 74 and 78 degrees by two observers 1km apart. how high is the plane?

OpenStudy (anonymous):

 angles on a line\[180-78=102\] angle opposite 1km 180 - 78 -102 = 4 find \[\frac{ \sin4 }{ 1 }= \frac{ \sin102 }{ ? }\] x= 15.9 fine height\[h = \sin74(15.9)\] \[=14.6\]